OpenStudy (anonymous):

There are 190mg of sodium in 27.0 g serving of Wheaties. Every morning Hector makes 75.0 g of a serving. How much sodium does Hector intake with his breakfast every morning? Show work

Parth (parthkohli):

Set up a proportion here. \(\Large \color{MidnightBlue}{\Rightarrow 190:27=x:75 }\) Since they have a direct relationship.

OpenStudy (kropot72):

Amount of sodium in a 1g serving is 190/27mg Amount of sodium in a 75 g serving is\[\frac{190}{27}\times 75\]

OpenStudy (anonymous):

10.66 g serving?

OpenStudy (anonymous):

wait no.

OpenStudy (anonymous):

When I set it up i got 190:27=x:75 27x/27=1330/27 =49.259 g of sodium

OpenStudy (kropot72):

\[\frac{190}{27}\times 75=527.78mg\]

OpenStudy (anonymous):

oh ok thank you

OpenStudy (kropot72):

You're welcome :)